OpenStudy (anonymous):

which values of x and y will make the expression 5(x-y)2 equal to 20?

jimthompson5910 (jim_thompson5910):

hint: let z = x-y replace "x-y" with z, then solve for z

OpenStudy (anonymous):

\[5(x-y)^{2} equal \to 20\]

OpenStudy (anonymous):

so is x-y/z

jimthompson5910 (jim_thompson5910):

5(x-y)^2 = 20 5z^2 = 20 z^2 = 20/5 z^2 = 4 z = ???

OpenStudy (anonymous):

z=8

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

z^2 means z squared

OpenStudy (anonymous):

i don't know

jimthompson5910 (jim_thompson5910):

if z^2 = 4, then z = 2 or z = -2 so if z = x-y, then x-y = 2 or x-y = -2

jimthompson5910 (jim_thompson5910):

from there, you can solve each equation for y to get it in terms of x